# Positive Semidefinite Matrix Supermartingales

### v4

@@ B.3 Proofs for Backward Submartingales
To see that there exists an $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ such that $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha)$, one can take $\mathbf{X}=\mathbf{A}$ as $\mathrm{tr}(\mathbf{A})\geqslant d^{2}/\alpha>d/\alpha$. To see that there exists an $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, suppose $\lambda_{\min}(\mathbf{A})=d/\alpha-2\delta$ where $\delta>0$. Take $\mathbf{X}$ to be a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $d/\alpha-\delta,\delta/(d-1),\dots,\delta/(d-1)$. ∎
@@
First, when $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}(\mathbf{X})\leqslant d/\alpha^{\prime}\leqslant 1/\lambda_{\max}(\mathbf{A}^{-1})=\lambda_{\min}(\mathbf{A})$ so $\mathbf{X}\in\mathcal{C}(\mathbf{A})$. Second, when $\alpha^{\prime}<\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\delta=(d/\alpha^{\prime}-\lambda_{\min}(\mathbf{A}))/2$ and take $\mathbf{X}$ as a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $\lambda_{\min}(\mathbf{A})+\delta,\delta/(d-1),\dots,\delta/(d-1)$ to see that $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ but $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$. ∎
@@
Complete the orthonormal basis $\{\mathbf{v}_{k}\}_{1\leqslant k\leqslant d}$ where $\mathbf{v}_{1}=\mathbf{v}$ and let $\mathbf{A}=\alpha^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}\delta\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for $\delta=\frac{\alpha^{\prime}-\alpha}{d-1}$. ∎
@@
$\mathbb{E}(L_{n}|\mathcal{F}_{n-1})=$ $\mathbb{E}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{n-1}\mathbf{Z}_{i}-\sum_{i=1}^{n-1}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})-\mathbf{C}_{n}^{\prime}+\log\mathrm{e}^{\mathbf{Z}_{n}-\mathbf{C}_{n}}\right)\middle|\mathcal{F}_{n-1}\right)$ (129)
@@
Appendix C Miscellaneous Technical Results
C.1 Matrix MGF Bounds
@@
Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
@@
The uni-Gaussian MGF bound is proved by Tropp [2015, Lemma 4.6.2] via a Taylor expansion argument. For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a scalar inequality similar to the Hoeffding’s inequality plus the transfer rule (see Section B.2). Bennett [Tropp, 2012, Lemma 6.7] and Bernstein [Tropp, 2012, Lemma 6.8] follow from a similar reasoning applied to the respective series expansion. ∎
C.2 A Matrix Doob’s $L^{p}$ Inequality
@@
Proposition C.3 (Doob-type Nuclear $L^{p}$ Inequality).
@@
$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{1}^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|^{p}_{1}=\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\|\mathbf{M}_{N}\|\cdot\operatorname{r}(\mathbf{M}_{N}))^{p}.$ (140)
@@
The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\|\mathbf{M}_{n}\|_{1}\}$. The first $\leqslant$ and the final $=$ follow from straightward relations between the spectral and nuclear norms, and the effective rank. ∎
C.3 Non-symmetric, Non-square Matrices
@@
Example C.4.
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds $\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable. For any predictable positive scalars $\{\gamma_{n}\}$, let

### v6

@@ 8.3 Proofs for Backward Submartingales
To see that there exists an $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ such that $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha)$, one can take $\mathbf{X}=\mathbf{A}$ as $\mathrm{tr}(\mathbf{A})\geqslant d^{2}/\alpha>d/\alpha$. To see that there exists an $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, suppose $\lambda_{\min}(\mathbf{A})=d/\alpha-2\delta$ where $\delta>0$. Take $\mathbf{X}$ to be a matrix with the same eigenvectors as $\mathbf{A}$ and eigenvalues $d/\alpha-\delta,\delta/(d-1),\dots,\delta/(d-1)$. ∎
@@
First, when $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}(\mathbf{X})\leqslant d/\alpha^{\prime}\leqslant 1/\lambda_{\max}(\mathbf{A}^{-1})=\lambda_{\min}(\mathbf{A})$ so $\mathbf{X}\in\mathcal{C}(\mathbf{A})$. Second, when $\alpha^{\prime}<\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\delta=(d/\alpha^{\prime}-\lambda_{\min}(\mathbf{A}))/2$ and take $\mathbf{X}$ as a matrix with the same eigenvectors as $\mathbf{A}$ and eigenvalues $\lambda_{\min}(\mathbf{A})+\delta,\delta/(d-1),\dots,\delta/(d-1)$ to see that $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ but $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$. ∎
@@
Complete the orthonormal basis $\{\mathbf{v}_{k}\}_{1\leqslant k\leqslant d}$ where $\mathbf{v}_{1}=\mathbf{v}$ and let $\mathbf{A}=\alpha^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}\delta\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for $\delta={\frac{d-1}{\alpha^{\prime}-\alpha}}$. ∎
@@
$\mathbb{E}(L_{n}|\mathcal{F}_{n-1})=\$$\mathbb{E}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{n-1}\mathbf{Z}_{i}-\sum_{i=1}^{n-1}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})-\mathbf{C}_{n}^{\prime}+\log\mathrm{e}^{\mathbf{Z}_{n}-\mathbf{C}_{n}}\right)\middle|\mathcal{F}_{n-1}\right)$ (133)
@@
9 Miscellaneous Technical Results
9.1 Matrix MGF Bounds
@@
Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
@@
The uni-Gaussian MGF bound is proved by Tropp [2015, Lemma 4.6.2] via a Taylor expansion argument. For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a scalar inequality similar to the Hoeffding’s inequality plus the transfer rule (see Section 8.2). Bennett [Tropp, 2012, Lemma 6.7] and Bernstein [Tropp, 2012, Lemma 6.8] follow from a similar reasoning applied to the respective series expansion. ∎
9.2 A Matrix Doob’s $L^{p}$ Inequality
@@
Proposition 9.3 (Doob-type Nuclear $L^{p}$ Inequality).
@@
$\scriptsize\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{1}^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|^{p}_{1}=\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\|\mathbf{M}_{N}\|\cdot\operatorname{r}(\mathbf{M}_{N}))^{p}.$ (144)
@@
The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\|\mathbf{M}_{n}\|_{1}\}$. The first $\leqslant$ and the final == follow from straightward relations between the spectral and nuclear norms, and the effective rank. ∎
9.3 Non-symmetric, Non-square Matrices
@@
Example 9.4.
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds
$\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable. For any predictable positive scalars $\{\gamma_{n}\}$, let